# MagicClay: Sculpting Meshes With Generative Neural Fields

### v1

@@ 3.1. Hybrid Representation
Our hybrid representation consists of a surface (a mesh), a volume (an SDF), and a shared appearance network encoding RGB colors for an input 3D coordinate. Both the surface and the volumetric representations can be differentiably rendered, leveraging the shared appearance network to output images with color, normals, and opacity channels. We now detail the three elements of our hybrid shape representation.
@@
We represent the surface of the shape as a 2-manifold triangular mesh. Mesh topology, or sampling resolution, is locally adapted according to the SDF (see Sec. 3.4 for details). We encode colors for the mesh using an auxiliary appearance network, derived from the SDF itself (see below). We found this approach simpler and more natural than traditional mesh coloring techniques; Using per-vertex colors is sensitive to triangulation, and would require a large number of vertices to match the resolution of the SDF. Using a texture image requires a complex UV parameterization, usually done a priori on a fixed shape. In addition, our surface is continuously optimized and undergoes through topological changes, re-tessellation, and large-scale deformations, which makes it computationally infeasible to apply traditional UV parameterization techniques during this optimization.
Instead, our hybrid approach offers a simpler approach to shape coloring. To apply the colors from the appearance network to the mesh, we propose to adaptively subdivide each face of the base mesh according to triangle area. Since we only use these subdivided triangles to represent colors they do not have to form a connected mesh unlike traditional subdivision techniques. Thus, we employ MeshColors scheme that was originally proposed for UV-less texturing [Yuksel et al., 2010], and has an efficient GPU implementation. In the inset we illustrate the example subdivision, note how sub-triangles on two adjacent faces do not share the vertices along the
edge. During rendering we assign a color to each sub-triangle, by using the coordinate of the three associated supersampled vertices to sample the appearance network.
@@
Our volumetric shape representation is chosen off-the-shelf, and conceptually serves as a regularization guiding the mesh evolution using existing state-of-the-art text2shape tools. We use a neural SDF, or an implicit continuous scalar field that can be sampled anywhere in $\mathbf{R}^{3}$, returning a signed shortest distance to the surface (negative on the inside, positive on the outside). We encode the SDF using a multiresolution hash encoding of features defined over a grid which are then mapped to distance value by a small MLP, following instant-NGP [Müller et al., 2022]. As in the mesh case, the shared appearance network is sampled to obtain colors during rendering.
@@ 3.2. Hybrid Shape Guidance
In essence, our shape optimization is based on Score-Distillation Sampling (SDS) to distill gradients from a text prompt. The primary motivation to maintain an SDF representation in addition to the mesh is because SDFs are more robust noisy guidance, which is an inherent property of the multi-view SDS approach (see Figure 5). We thus choose to inject the text guidance only to the auxiliary SDF representation, and propagate the changes to the mesh via the consistency losses (Sec. 3.3) and the topology updates (Sec.  3.4).
@@
This is achieved thanks to the consistency between the mesh and SDF representations throughout the optimization. We can significantly reduce the typical 512 samples per ray necessary for rendering the SDF by using the intersection of the ray with the mesh representation (efficiently calculated by the differentiable mesh renderer). Using the intersection as the center of a small spread of samples (typically 3), this allows for high resolution renders of the SDF (i.e. 512x512 and larger), which are otherwise memory prohibitive. The idea of leveraging the surface to reduce the number of network queries per ray emerged in concurrent works, namely Adaptive Shell [Wang et al., 2023b] and HybridNerf [Turki et al., 2023], which shows its generality and success in other settings than ours.
Using this strategy, we render the SDF in 512x512 and apply the VSD loss of those high-res renderings. We also aply regular SDS on lower-res 128x128 renderings by regular VolSDF as we find that this improves the results slightly.
@@ 3.3. Representation Priors
In addition to the text guidance, we apply representation-specific regularizations and consistency losses that keep both representations in sync.
@@
The SDF and the mesh are consistent if their images are in 1 to 1 correspondences from any camera angle. We thus supervise the L2 difference between their RGB renderings, normal maps and opacity maps. If the renderings are made at different resolution, we downsize to the lower resolution before computing the L2 loss.
@@
To localize changes to the user-selected area we first fix the mesh vertices in all non-selected regions during optimization by zeroing out gradients outside of user selection. While localization is harder to achieve for SDF, we add a sampling-based freeze loss, which favors regions around fixed vertices to remain unchanged:
@@
While it is harder to regularize the surface of an implicit function to be smooth, the explicit representation of the mesh allows to easily define a smoothness term using the Laplacian of the mesh, defined for each vertex:
(2) $\delta(x_{i})=x_{i}-\frac{\Sigma_{j}{x_{j}}}{N},$
where $x_{j}$ are neighbors of $x_{i}$. The Laplacian vector encodes local geometry changes, with a smooth mesh is defined by low Laplacian vectors, and we use a global smoothness loss:
(3) $L_{\text{smooth}}=\Sigma_{i}{||\delta_{i}||}.$
@@
To encourage the implicit function to learn a valid SDF representation we use the Eikonal term as a loss. The SDF $s$ is valid if and only if the loss in Eqn 4 is 00:
@@
Inspired by TextMesh [Tsalicoglou et al., 2023], we also binarize the SDF opacity and apply a Binary Cross Entropy loss to encourage discrete 0 or 1 values. To penalize badly oriented normals of the implicit surface, we apply an L2 penalty to the dot product between the normal and the camera direction if it is negative.
@@ 3.4. Updating the mesh topology
To maintain consistency between the mesh and SDF, it is necessary to perform local topology edits on the mesh in that increase or decrease mesh resolution where required. Continuous Remeshing [Palfinger, 2022] pioneered such a local topology update approach, by using the Adam optimizer state as a signal. While this approach works well in a multi-view reconstruction scenario, where the images are sharp, and the camera parameters known, the noise involved in SDS makes the gradients, and by extension the Adam state, very noisy and an unstable signal to trigger those operations. We turn to another work, ROAR [Barda et al., 2023], particularly well-tailored to our hydrid representation. Within this framework, we use the SDF as the signal to trigger mesh triangle splits.
@@
This projection results in a piece-wise linear surface that approximate the implicit surface closest to the initial triangle. The decision to split this triangle is based on the curvature score of this piece of projected surface. If the surface is very curved, then the triangle is split using $\sqrt{3}$-subdivision [Kobbelt, 2000]. Similarly, each edge is assigned a score based on the quadratic distance of its vertices to all the planes in the 1-ring of the edge, which intuitively represents how important is the edge to the geometry. If the score is low, then the edge can be collapsed with Qslim [Garland and Heckbert, 2023].
We refer the interested reader to the ROAR paper [Barda et al., 2023] for more details, but the important point to note that ROAR offers a principled way to perform edge collapses and face splits in the sense that each iteration of ROAR strictly decrease an energy - the difference between the highest face score and the lowest edge score. It thus exhibits a convergence behavior after enough iterations. We also note that manifoldness is guaranteed to be preserved throughout the iterations.
@@ 4. Experiments
We implement our pipeline in Threestudio [Guo et al., 2023], and use the implementations of other methods provided in the framework for comparisons with Stable Diffusion v1.5 as the backbone diffusion model. All experiments presented in this work were executed on a single A100-40GB GPU.
In the rest of this section we compare our representation to prior work on text-conditioned 3D generation (Sec. 4.1), demonstrate its utility in a mesh sculpting application (Sec. 4.2), and compare to a text-driven mesh deformation baseline (Sec. 4.3). We then provide a simple illustrative experiment to motivate the hybrid representation when using SDS guidance with noisy gradients (Sec 4.4), and finally ablate our method (Sec 4.5).
Figure 3. Comparison on text-to-3D from scratch. We compare the quality of the triangular meshes extracted from various state-of-the-art generative methods: Fantasia3d [Chen et al., 2023], ProlificDreamer [Wang et al., 2023a] and TextMesh [Tsalicoglou et al., 2023]. While all methods produce realistic RGB renderings, only our hybrid representation generates smooth geometry.
@@ 4.1. Comparison with Generative Methods
Since MagicClay is a modeling tool, we are primarily interested in evaluating the quality of the geometry and thus focus on mesh renderings without texture. Note that existing 3D generative techniques are not designed to edit a part of an existing mesh, and thus we compare performance of our hybrid approach on the task of text-to-3D generation. We compare against three recent approaches: Fantasia3d [Chen et al., 2023], ProlificDreamer [Wang et al., 2023a] and TextMesh [Tsalicoglou et al., 2023]. The representation used in TextMesh [Tsalicoglou et al., 2023] is an SDF, so we run marching cube on their output to evaluate the quality of the mesh. ProlificDreamer uses a two stage approach: first a NeRF-base generation, then a refinement using an explicit representation. Fantasia3d uses DMtet [Shen et al., 2021] to extract a mesh from an SDF.
We present our results in Figure 3. Even though each approach generates a representation that can produce realistic RGB renders, extracted meshes often exhibit significant surface artifacts, which make them hardly recognizable without texture (see “Chow Chow puppy” by ProlificDreamer or “Croissant” by TextMesh). By comparison, our geometries are recognizable and smooth thanks to the fact that our hybrid representation enables an explicit regularization of the surface. This validates that MagicClay successfully bridges the generative capabilities of implicit radiance fields with the surface-level controls of meshes.
@@ 4.2. Mesh Sculpting
We further demonstrate that our method can be used to enable novel iterative 3D sculpting workflows. Starting with an initial mesh, an artist can select region of interest (and optionally sculpt a coarse adjustment to that region) along with a textual prompt describing the desired updated shape. MagicClay generates a modified mesh, which could be iteratively refined with new elements. Note that hybrid representation is essential to this application. First, selecting and adjusting the region of interest is easily accomplished using standard mesh editing tools [Blender, 2024]. Second, the generated result tends to be more expressive with additional SDF representation guidance. Third, using the mesh allows us to keep non-selected surface regions intact by zeroing out their deformation gradients, which guarantees that the change will only affect the user-selected region. Refer to Figures 19 and supplemental video for example results. MagicClay generates a high quality edits, that matche the rough local edit and adhere to the user’s text prompt.
Figure 4. Comparisons to TextDeformer [Gao et al., 2023]. Given the user-modified input mesh, we either run TextDeformer or our method to modify it towards the target prompt.
4.3. Comparison with TextDeformer
Our method is the first to tackle the problem of interactive, localized mesh sculpting via text-based prompts. A naive alternative would be to simply use the existing text-driven mesh deformation technique [Gao et al., 2023] on the user-modified mesh, still aiming to achieve desired changes in the geometry. In Figure 4 we compare our method to this baseline. Note how our method is able to add geometrically complex large-scale details due to guidance from SDF and topological updates. Our method’s changes are also restricted only to user-modified region, and do not lead to large-scale deformations in the other parts of the input.
@@ 4.4. Analysis of Mesh and SDF robustness to noise
Figure 5. Mesh and SDF robustness to noisy gradients. We optimize a mesh, an SDF our our hybrid representation with multi-view reconstruction losses after applying various noise levels to the ground truth renderings. We report the L2 reprojection error against the ground truth renders. The SDF exhibit more robustness than the mesh to high noise regime, and our hybrid outperforms both.
We now illustrate the motivation for our hybrid representation by a simple controlled experiment, where we aim to reconstruct a fixed 3D target with different levels of noise in the guidance. We formulate it as a reconstruction problem to have a clear ground truth and eliminate ambiguity arising due to text-based objectives. Even though we use synthetic noise, we expect these findings to apply in an SDS setting, where gradients are also noisy due to random noising step performed at each SDS iteration [Poole et al., 2022].
Given multi-view renderings of a fixed (true) 3D model, we add uncorrelated per-pixel Gaussian noise to each image, and compute L2 pixel-wise loss to guide our shape representation towards the target. As we increase the noise level (by increasing standard deviation) we find that different representations are more prone to errors in reconstructing the target. We use L2 re-projection error with respect to the ground truth shape as our evaluation metric, and compare vanilla Mesh, SDF representations to our hybrid approach (using our mesh rendering), and show results in Figure 5. Note that as the amount of noise increases, the quality of the mesh reconstruction degrades sharply. At the highest noise regimes (standard deviation of 5), the mesh reconstruction degenerates to a blob, while the SDF reconstruction is still recognizable despite surface irregularities. Importantly, the hybrid representation performs better than both individual representation at all levels of noise, and the benefits are the strongest at higher noise level.

### v2

@@ 3.1. Hybrid Representation
Our hybrid representation consists of a surface (a mesh), a volume (an SDF), and a shared appearance network encoding RGB colors for an input 3D coordinate. Both the surface and the volumetric representations can be differentiably rendered, leveraging the shared appearance network to output images with color, normals, and opacity channels. We now detail these three elements.
@@
We represent the surface of the shape as a 2-manifold triangular mesh. Mesh topology, or sampling resolution, is locally adapted according to the SDF (see Sec. 3.4 for details). We encode colors for the editable areas of the mesh using an auxiliary appearance network, derived from the SDF itself (see below). We found this approach simpler and more natural than traditional mesh coloring techniques; Using per-vertex colors is sensitive to triangulation, and would require a large number of vertices to match the resolution of the SDF. Using a texture image requires a complex UV parameterization, usually done a priori on a fixed shape. In addition, our surface is continuously optimized and undergoes topological changes, re-tessellation, and large-scale deformations, which makes it computationally infeasible to apply traditional UV parameterization techniques during this optimization.
Instead, our hybrid approach offers a simpler approach to shape coloring. To apply the colors from the appearance network to the mesh, we propose to adaptively subdivide each face of the base mesh according to the triangle area.
Since we only use these subdivided triangles to represent colors, they do not have to form a connected mesh, unlike traditional subdivision techniques. Thus, we employ the MeshColors scheme that was initially proposed for UV-less texturing [Yuksel et al., 2010] and has an efficient GPU implementation.
In the inset, we illustrate the example subdivision; note how sub-triangles on two adjacent faces do not share the vertices along the
edge. During rendering, we assign a color to each sub-triangle by interpolating the colors at its three vertices. These are obtained by querying the appearance network at the super-sampled vertices location.. We find that this approach strikes a balance between a compact low-poly representation for flat parts of the generated surface, while still allowing for high-frequency textures.
@@
Our volumetric shape representation is chosen off-the-shelf, and conceptually serves as a regularization guiding the mesh changes using existing state-of-the-art text2shape tools. We use a continuous scalar field that can be sampled anywhere in $\mathbf{R}^{3}$, returning a signed shortest distance to the surface (negative on the inside, positive on the outside). We encode the SDF using a multiresolution hash encoding of features defined over a grid which are then mapped to distance value by a small MLP, following instant-NGP [Müller et al., 2022]. As in the mesh case, the shared appearance network is sampled to obtain colors during rendering.
@@ 3.2. Hybrid Shape Guidance
Our shape optimization is based on Score-Distillation Sampling (SDS) to distill gradients from a text prompt. The primary motivation to maintain an SDF representation in addition to the mesh is because SDFs are more robust to noisy guidance, which is an inherent property of the multi-view SDS approach (see Figure 3). We thus choose to inject the text guidance only to the auxiliary SDF representation and propagate the changes to the mesh via the consistency losses (Sec. 3.3) and the topology updates (Sec.  3.4).
@@
This is achieved thanks to the consistency between the mesh and SDF representations throughout the optimization.
We can significantly reduce the typical 512 samples per ray necessary for rendering the SDF by using the intersection of the ray with the mesh representation (efficiently calculated by the differentiable mesh renderer). Using the intersection as the center of a small spread of samples (we use just 3 per ray), allows for high-resolution renders of the SDF (i.e., 512x512 and larger), which are otherwise memory prohibitive. The idea of leveraging the surface to reduce the number of network queries per ray emerged in concurrent works, namely Adaptive Shell [Wang et al., 2023b] and HybridNerf [Turki et al., 2024], which shows its generality and success in other settings than ours.
Using this strategy, we render the SDF in 512x512 and apply the VSD loss of those high-res renderings. We also apply regular SDS on lower-res 128x128 renderings by regular VolSDF as we find that this improves the results slightly.
@@ 3.3. Representation Priors
We apply representation-specific regularizations and consistency losses that keep both representations in sync.
@@
The SDF and the mesh are consistent if their images are in 1 to 1 correspondences from any camera angle. We thus supervise the L2 difference between their RGB renderings and normal and opacity maps. If the renderings are made at different resolutions, we downsize to the lower resolution before the L2 loss.
@@
To localize changes to the user-selected area, we first fix the mesh vertices in all non-selected regions during optimization by zeroing out gradients outside of user selection. While localization is harder to achieve for SDF, we add a sampling-based freeze loss, which favors regions around fixed vertices to remain unchanged:
@@
While it is harder to regularize the surface of an implicit function to be smooth, the explicit representation of the mesh allows to easily define a smoothness term using the Laplacian of the mesh, inspired by [Kanazawa et al., 2018], defined for each vertex:
(2) $\delta(x_{i})=x_{i}-\frac{\Sigma_{j=1}^{N}{x_{j}}}{N},$
where $x_{j}$ are the N neighbors of $x_{i}$. The Laplacian vector encodes local curvature changes: a smooth mesh is defined by low Laplacian vectors. To encourage smoothness, we use a global loss:
(3) $L_{\text{smooth}}=\Sigma_{i}{||\delta(x_{i})||}.$
We opt for the uniform Laplacian instead of the cotan Laplacian, as the latter is more sensitive to ill-conditioned triangles, which may appear during the dynamic mesh updates.
@@
To encourage the implicit function to learn a valid SDF representation, we use the Eikonal term as a loss  [Gropp et al., 2020]. The SDF $s$ is valid if and only if the loss in Eqn 4 is 00:
@@
Inspired by TextMesh [Tsalicoglou et al., 2024], we also binarize the SDF opacity and apply a Binary Cross Entropy loss to encourage discrete 0 or 1 values. To penalize badly oriented normals of the implicit surface, we apply an L2 penalty to the dot product between the normal and the camera direction if it is negative.
@@ 3.4. Updating the Mesh Topology
To maintain consistency between the mesh and SDF, it is necessary to perform local topology edits on the mesh in that increase or decrease mesh resolution where required. Continuous Remeshing [Palfinger, 2022] pioneered such a local topology update approach by using the Adam optimizer state as a signal. While this approach works well in a multi-view reconstruction scenario, where the images are sharp and the camera parameters known, the noise involved in SDS makes the gradients, and by extension the Adam state, very noisy and unstable signal to trigger those operations. We turn to another work, ROAR [Barda et al., 2023], particularly well-tailored to our hydrid representation. Within this framework, we use the SDF as the signal to trigger mesh triangle splits.
@@
This projection results in a piece-wise linear surface that approximates the implicit surface closest to the initial triangle. The decision to split this triangle is based on the curvature score of this piece of projected surface. If the surface is very curved, then the triangle is split using $\sqrt{3}$-subdivision [Kobbelt, 2000]. Similarly, each edge is assigned a score based on the quadratic distance of its vertices to all the planes in the 1-ring of the edge, which intuitively represents how important the edge is to the geometry. If the score is low, then the edge can be collapsed with Qslim [Garland and Heckbert, 2023].
We refer the interested reader to the ROAR paper [Barda et al., 2023] for more details, but the important point to note is that ROAR offers a principled way to perform edge collapses and face splits in the sense that each iteration of ROAR strictly decreases an energy - the difference between the highest face score and the lowest edge score. It thus exhibits a convergence behavior after enough iterations. We also note that manifoldness is guaranteed to be preserved throughout the iterations.
@@ 4. Experiments
Appearance Geometry
Average Implicit Implicit Implicit Mesh no texture
CLIP Score $\uparrow$ T2I type RGB Normal Normal
ProlificDreamer SD2.1 NeRF 22.1 21.2 20.1
HIFA SD2.1 NeRF 23.2 23.1 22.0
MVDreams SD2.1 NeRF 25.9 25.4 24.1
Fantasia3D SD1.5 SDF - 23.9 23.9
TextMesh DF SDF 25.7 24.9 23.6
Ours DF SDF 26.2 26.1 24.8
1 No high-res DF SDF 25.8 25.2 23.1
2 No low-res DF SDF 22.3 21.9 21.1
Table 1. Quantitative comparison on text-to-3D from scratch. To validate our hybrid representation, we compare MagicClay with five state-of the-art method for unconditional generation. We report the CLIP Score, i.e. the cosine similarity (scaled by 100) between prompt embeddings and 25 normal renders of the NeRFs/SDFs and texture-less meshes, taken in a equidistant circular pattern around the generated object, on a benchmark of 20 prompts, listed in the supplementary. We apply marching cubes to extract the meshes for NeRF-based methods, and isosurface extraction for SDF-based methods, except ours, as our dual representation already includes a mesh. For each method, we report the Text-to-Image model (T2I) used as a backbone : DeepFloyd (DF) or Stable-Diffusion (SD). We Note that Fantasia3D can not be run with DF, as it does not support diffusion in pixel-space. We validate that MagicClay has the highest Clip score for texture-less meshes, which can be verified qualitatively in Figure 4.
We implement our pipeline in Threestudio [Guo et al., 2023], and use DeepFloyd  [StabilityAI, 2023] as the backbone diffusion model. All experiments were executed on a single A100-40GB GPU. We provide further details, hyper-parameters, and additional examples of the texture maps generated by our method and the animation transfer application in the supplementary material.
In the rest of this section, we compare our representation to prior work on text-conditioned 3D generation (Sec. 4.1), demonstrate its utility in a mesh sculpting application (Sec. 4.2), and compare to a text-driven mesh deformation baseline (Sec. 4.3). We then provide a simple illustrative experiment to motivate the hybrid representation, and finally ablate our method (Sec. 4.5).
@@ 4.1. Comparison with Generative Methods
Since MagicClay is a modeling tool, we are primarily interested in evaluating the quality of the geometry and thus focus on mesh renderings without texture. Note that most existing 3D generative techniques are not designed to edit a part of an existing mesh, and therefore we start by comparing the performance of our hybrid approach for unconditional text-to-3D generation, seeking to establish the ability of our method to generate higher-quality meshes. We compare against five recent approaches: HIFA [Zhu and Zhuang, 2024], MVDream [Shi et al., 2024], Fantasia3D [Chen et al., 2023], ProlificDreamer [Wang et al., 2023a] and TextMesh [Tsalicoglou et al., 2024]. We run the open-source implementation of all methods to produce results. We use Marching Cube to extract a mesh from NeRF-based methods, and isosurface extraction for SDF-based methods. We do not compare against Instant3D [Li et al., 2024] because they do not provide an implementation as of writing this paper, but we note their ability to generate high-quality meshes. We measure the quality of the generated objects using an average clip score between the prompt and multi-view renderings. A complete description of our metric and the prompts used are given in the supplementary materials. We present our results in Figure 4 and Table 1. Extracted meshes often exhibit significant surface artifacts, which make them hardly recognizable without texture (see “Chow Chow puppy” by ProlificDreamer or “Croissant” by TextMesh). By comparison, our geometries are recognizable and smooth thanks to our hybrid representation enabling an explicit regularization of the surface. We further verify that, among competing approaches, MagicClay achieves the highest Clip Score between the prompts and texture-less renderings of the mesh. This shows that MagicClay achieves semantically meaningful reconstruction via geometry and not only texture, and thus successfully bridges the generative capabilities of implicit radiance fields with the surface-level controls of meshes.
Our hybrid representation is orthogonal to the choice of backbone text-to-image models. Fig 9 shows 3D models generated by MagicClay with SD 1.5, SD 2.1, and DeepFloyd, showing that the backbone model influences the quality of the results. This drives two conclusions : (1) improvements in the backbone translate into improvements for 3D generation. (2) In addition to the representation, an important differentiating factor for each method in Table 1 is the backbone, which we report.
@@ 4.2. Mesh Sculpting
MagicClay generates a modified mesh, which could be iteratively refined with new elements. Note that hybrid representation is essential to this application. Selecting the region of interest is easily accomplished by using standard mesh editing tools [Blender, 2024]. Using the mesh allows us to keep non-selected surface regions intact by zeroing out their deformation gradients, which guarantees that the change will only affect the user-selected region. Because they are preserved, mesh properties can be transferred in the non-edited region, including texture, tessellation, and rigging parameters. Refer to Figures 1 for example results. MagicClay generates high-quality edits that match the rough local edit and adhere to the user’s text prompt.
4.3. Comparison with Contemporary 3D Editing Pipelines
A naive alternative to localized mesh sculpting would be to use the existing text-driven mesh deformation technique [Gao et al., 2023] and to zero-out the deformation field outside of the editing region, which is possible because the representation is explicit. We further compare with Latent-Nerf [Metzer et al., 2023] and Vox-E [Sella et al., 2023], but these approaches cannot guarantee a local edit because they are based on soft attention. We discuss InstructNeRF2NeRF [Haque et al., 2023] and DreamEditor [Zhuang et al., 2023] in supplementary
In Figure 8, we compare our method to these three baselines. Note how our method is able to add geometrically complex large-scale details due to guidance from SDF and topological updates. Our method’s changes are also restricted only to user-selected regions and do not lead to any changes in the other parts of the input.
@@ 4.4. Analysis of Mesh and SDF Robustness to Noise
Figure 3. Mesh and SDF robustness to noisy gradients. We optimize a mesh, an SDF and our hybrid representation with multi-view reconstruction losses after applying various noise levels to the ground truth renderings. We report the L2 reprojection error against novel-views ground truth renders. The SDF exhibits more robustness than the mesh to the high noise regime. We show the results for both the mesh (hybrid-mesh) and SDF (hybrid-SDF) in our hybrid representation. The hybrid-mesh significantly outperforms the mesh only baseline in the high noise regime.
We now motivate the use of our hybrid representation by a simple controlled experiment, where we aim to reconstruct a fixed 3D target with different levels of noise in the guidance. Even though we use synthetic noise, we expect these findings to apply in an SDS setting, where gradients are noisy due to the noising step performed at each SDS iteration [Poole et al., 2023].
Given multi-view renderings of a fixed ground truth 3D model, we add uncorrelated per-pixel Gaussian noise to each image and compute L2 pixel-wise loss to guide our shape representation towards the target. A complete overview of the experiment setup is given in the supplementary materials. As we increase the noise level (by increasing standard deviation) we find that different representations are more prone to errors in reconstructing the target. We use L2 re-projection error with respect to the ground truth shape on novel views as our evaluation metric, and compare vanilla Mesh, SDF representations to our hybrid approach (using both SDF and mesh renderings), and show results in Figure 3. At the highest noise regimes (standard deviation of 5), the mesh reconstruction degenerates to a blob, while the SDF reconstruction is still recognizable despite surface irregularities. Importantly, the hybrid representation performs better than both individual representations at all levels of noise, and the benefits are the strongest at higher noise levels. The hybrid-SDF outperforms the SDF baseline because the mesh acts as a regularizer for more 3D consistency. The L2 reprojection loss of the hybrid-mesh is lower than the hybrid-SDF because the SDF tends to exhibit a “haze” artifacts throughout the entire image, including the background pixels, while the mesh rendering can not create this effect in the background, serving as a sort of regularization on its own. Finally, we note that all curves are not strictly increasing, which we attribute to the fact that we test on novel views: some amount of noise may prevent overfitting to the training views.
Note that this experiment is not designed to claim superiority of the hybrid representation over the SDF, but rather to show its robustness to noise.